OpenStudy (chris215):

Find the derivative of the function f(x) = cos^2(x) + tan^2(x)

OpenStudy (chris215):

I got 2cos(x)sin(x) + sec2(x)

OpenStudy (michele_laino):

please keep in mind the the first derivative of \(\cos x\) is \(-\sin x\)

OpenStudy (michele_laino):

namely: \[\frac{d}{{dx}}{\left( {\cos x} \right)^2} = 2\cos x\left( {\frac{d}{{dx}}\cos x} \right) = - 2\cos x\sin x\]

OpenStudy (chris215):

ok right

OpenStudy (michele_laino):

:)

OpenStudy (chris215):

I'm confused what do that tell me?

OpenStudy (chris215):

wait now im get 2tan(x)sec^2(x) - 2sin(x)cos(x)